OpenStudy (anonymous):

please help me? p - 4 = -9 + p i will give a medal for the best response. :)

pooja195 (pooja195):

Are we solving for p?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well what do you think it is?

pooja195 (pooja195):

Add 4 to both sides -9+4=?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

i got to p - 5 = p but then i didnt know what to do i think i did it wrong

OpenStudy (anonymous):

No close

OpenStudy (anonymous):

do -9-4 on a caulculator

pooja195 (pooja195):

p−4=−9+p p+−4=−9+p p−4=p−9 Subtract p from both sides. p−4−p=p−9−p −4=−9 Add 4 to both sides. −4+4=−9+4 0=−5 When you see this it means that there is no solution #=# means infinte #= a different number means NO SOLUTION thereofre the answer would be no solution

OpenStudy (anonymous):

wow! thank you so much!!

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)